OpenStudy (anonymous):

If f(x) = |x|, for what value of x does f(x-2) = f(x+2)?

myininaya (myininaya):

|x-2|=|x+2| It is asking you to solve this.

Parth (parthkohli):

@CLester13 So, do you have a better way to say, "f(x - 2) = f(x + 2)"?

Parth (parthkohli):

What myin said. ^

Parth (parthkohli):

If you think like a philosopher, |-something| = |something| right?

Parth (parthkohli):

Or, |-something| = |-something| or |something| = |something| Which leaves JUST ONE POSSIBILITY: x - 2 is the opposite of x + 2, right?

Parth (parthkohli):

@CLester13: Hullo.

Parth (parthkohli):

Okay, answer this thing when you come online: How do you represent this sentence - "the opposite of x + 2"?

OpenStudy (anonymous):

x-2

myininaya (myininaya):

I would square both sides of |x-2|=|x+2| so I wouldn't need to worry about the absolute value. So perhaps you know how to solve this: \[(x-2)^2=(x+2)^2\] Hint: Expand both sides. You will be in for a real treat.

Parth (parthkohli):

That's a really cool way. :)